For P=(5,2) and Q=(7,-11), Δx, Δy, and the slope of the line segment from P to Q. (use decimal value for fractional answer)
a. Δx = ___
b. Δy = ___
c. slope = ___

Found 3 solutions by josgarithmetic, Theo, Shin123:
Answer by josgarithmetic(39620)
Answer by Theo(13342)

delta x means change in the value of x.
delta y means change in the value of y.

you have 2 points.
they are (5,2) and (7,-11)

slope is equal to (y2-y1) / (x2-x1)

let the point (5,2) be equal to (x1,y1)
let the point (7,-11) be equal to (x2,y2)

slope is equal to (y2-y1) / (x2-x1) = (-11 - 2) / (7 - 5) = -13/2.

-13/2 is equivalent to -6.5 in decimal form.

change in the value of x is equal to (x2 - x1) = (7 - 5) = 2

change in the value of y is equal to (y2 - y1) = (-11 - 2) = -13.

your solutions are:

change in the value of x is equal to 2
change in the value of y is equal to -13
slope is equal to -6.5





Answer by Shin123(626)

Δx is the change in x and Δy is the change in y. Slope=(y2-y1)/(x2-x1) (-11-2)/(7-5) -13/2.
Δx=2
Δy=-13
Slope=-6.5